# Size Generalization of Graph Neural Networks on Biological Data: Insights and Practices from the Spectral Perspective

### v4

@@ Appendix I Additional Results on Real-world Datasets
Let $C$ be a finite set of node features, and let $G=(V,E)$ be a graph with node feature $c_{v}\in C$ for every node $v\in V$. We define the d-pattern of a node $v\in V$ for $d\geq 0$ recursively: For $d=0$, the 0-pattern is $c_{v}$. For $d>0$, the $d$-pattern of $v$ is $p=(p_{v},\{(p_{i_{1}},m_{i_{1}}),\ldots,(p_{i_{l}},m_{i_{l}})\})$ iff node $v$ has $(d-1)$-pattern $p_{v}$ and for every $j\in\{1,\ldots,l\}$ the number of neighbors of $v$ with $(d-1)$-patterns $p_{i_{j}}$ is exactly $m_{p_{i_{j}}}$. Here, $i$ is the number of distinct neighboring $d-1$ patterns of $v$.
@@
We can prove this statement through induction on $d$. The 1-pattern corresponds to the node’s degree. In any cycle graph, the degree of any node is the same, which equals 2. For induction, we assume the statement holds for $k$-pattern, i.e. the $k$-pattern of every node is $\mathcal{K}$. Then we aim to prove the statement for the $(k+1)$-pattern. Note that every node has two neighboring nodes. By the recursive update rule of $d$-pattern, the $(k+1)$-pattern of any node is $(\mathcal{K},\{(\mathcal{K},\mathcal{K})\})$ and thus the $(k+1)$-pattern of every node is identical.
@@
We can prove this statement through induction on $d$. The statement can be easily verified for $d=0$ or $1$. Suppose the statement holds for $d\leq k$ and we want to show that it holds for the $(k+1)$-pattern. Note that the statement naturally holds for node $0,4$ since their $(k+1)$-pattern is dependent on the $k$-pattern of the same neighbor nodes. WLOG it suffices to consider the $k+1$-pattern node 2 for the rest of the cases, where we can enumerate all the recursive update cases from Figure 6. It is clear to observe that in each hierarchy the dependent neighbor patterns are the same and thus the d-pattern is identical. Hence we prove this claim.

### v5

@@ Appendix I Additional Plots on Scaled Eigenvector
Figure 5. A simple counter-example of d-pattern. The d-pattern for each pair of nodes with the same id (number) in figures (a) and (b) is identical (proved in K.3). The figure follows the same convention as Figure 2 of Yehudai et al. (2021), where the left side represents the two non-isomorphic graphs with 7 nodes and each color represents a different feature. The right side illustrates the 0, 1, and 2-pattern of the node with id 2.
Figure 6. Illustration of recursive updates of node 2 in (a) and (b) from Figure 5. In each hierarchy, the d-pattern is the same.
@@
[…]
@@
[…]
@@
[…]
BBBP
BACE
PROTEINS
NCI1
Figure 7. Pairwise graph distance measured by the Wasserstein distance of eigenvalue distributions after breaking cycles and aligning cycle lengths on different datasets. Breaking cycles amplifies the correlation between eigenvalue distribution and graph size, while aligning cycle lengths reduces the correlation.